OpenStudy (anonymous):

How to integrate [x*cos((5x))^2]dx ?

OpenStudy (turingtest):

use\[\cos^2x=\frac12[1+\cos(2x)]\]and integrate by parts

OpenStudy (turingtest):

we are talking about \[\int x\cos^2(5x)dx\]right?

OpenStudy (anonymous):

yeah :)

OpenStudy (turingtest):

ok, follow my prescription above what part are you stuck on?

OpenStudy (anonymous):

Got the answer :) thanks for the help :)

OpenStudy (turingtest):

welcome :)